# Machine-generated text detection prevents language model collapse

### v4

@@ 6 Conclusion
This work investigates model collapse across three dimensions: model performance, data quality, and resemblance to human samples. Through our analysis we found that the extent of model collapse and the effect on the data distribution is influenced by the decoding strategy. Truncating can lead to peaked distributions and repetitive models while sampling directly can result in high perplexity and verbose outputs with low resemblance to human data. Using the decoding strategies that resulted in the most extreme collapse, we evaluated the partially synthetic scenario, where human data is mixed into the training data. We designed a novel method to mitigate model collapse based on resampling the training distribution using the predictions of a machine-generated text detector. We have validated our method on two popular LLMs (GPT-2 and SmolLM2) across a range of decoding strategies, showing that we can prevent model collapse in all cases. When there is an equal ratio of human to synthetic samples in the training pool, our method results in improved model performance compared to training only on the human data.
@@
The primary limitation of this study is the evaluation of relatively small LLMs ($117$M and $360$M parameters) compared to larger models, which often have billions of parameters, limiting the generalisability of these findings. However, recent work Dohmatob et al. (2024a) suggests that model collapse is even more pronounced in larger models, indicating that our proposed approach could be even more beneficial in such cases. Future work could extend this analysis to larger models to validate these findings.
Additionally, as in previous studies Shumailov et al. (2023); Dohmatob et al. (2024b), we assess LLMs exclusively in a fine-tuning setting rather than pre-training from scratch. While pre-training experiments could provide deeper insights, the computational cost and complexity of training large-scale models from the ground up make such an approach impractical in our case. Nevertheless, given that model collapse has been primarily evaluated in LLMs from a fine-tuning setting, the conclusions made in this work still align with the current body of research.
Another limitation is that our study focuses primarily on open-ended text generation tasks. While this is a crucial area for understanding model collapse, our findings may not fully generalise to other domains, such as structured prediction or code generation, where the impact of model collapse may manifest differently. Future work could explore whether our resampling method remains effective across these domains.
Acknowledgments
G. Drayson is funded by the EPSRC grant “AI Centre for Doctoral Training in Foundational Artificial Intelligence” (EP/S021566/1). V. Lampos would like to acknowledge the support from the EPSRC grant “Digital Health Hub for Antimicrobial Resistance” (EP/X031276/1).
@@ Acknowledgments
Appendix A Machine-generated text detection
@@ A.1 Dataset
We trained and evaluated the machine-generated text detectors on the MAGE dataset Li et al. (2024), which is based on documents from $10$ domains: opinion statements (CMV & Yelp reviews dataset), news articles (XSum & TLDR dataset), question answering (ELI5), story generation (Reddit WritingPrompts & ROC), commonsense reasoning (HellaSwag), knowledge illustration (SQuAD) and Scientific writing (SciGen). The authors sampled $1{,}000$ texts from each domain (apart from opinion statements and news articles with $804$ and $777$ samples respectively) and generated text using $27$ LLMs from $7$ model families, which include OpenAI, LLaMA, GLM, FLAN-T5, OPT, BigScience and EleutherAI. For each human-written sample in the dataset, they generate a machine-generated version by providing the first $30$ tokens of human-written text as context to the LLM. In addition, for the OpenAI models, they implemented two other prompt strategies for relevant domains: ‘topical’ prompts such as an argument or news title and ‘specified’ prompts which contain information about the domain source. This results in $33{,}000$ ($27{,}000\!+\!3\times 2\times 1{,}000$) machine-generated samples per source before processing and filtering. The authors split the dataset into train, validation and test splits in the ratio $80:10:10$. To mitigate data imbalance in the validation and test sets they sample additional human data from each data source. The resulting test set contains $28{,}741$ human and $28{,}078$ machine-generated samples ($49\%$ machine-generated). The training set, however, is $71\%$ machine-generated. The total dataset consists of $154{,}078$ human-written and $294{,}381$ machine-generated texts. In addition to the test set, we also evaluate our detector on their more challenging test set containing text from four unseen domains (CNN/DailyMail, DialogSum, PubMedQA, IMDB) and generated by an unseen model (GPT-4). This out-of-distribution test set contains $762$ human and $800$ machine-generated samples.
@@
Figure S1: Wikitext test set performance (perplexity and accuracy) for varying mixing coefficients ($\alpha$, $\beta$, $\gamma$) across generations 00 to $9$ using pure sampling decoding with GPT-2.
@@
A.2 Pre-trained models
A.3 Hyperparameters
Each model was fine-tuned for 5 epochs with an early stopping patience of 3 epochs which is based on the macro average F1 score on the validation set and with no warm-up. Models were evaluated on the validation set after every epoch which was used to select the best model for evaluation. Optimisation was performed using AdamW with a $\beta_{1}$, $\beta_{2}$, $\epsilon$, and weight decay of $0.9$, $0.98$, $10^{-6}$, and $10^{-2}$, respectively. These parameters were chosen based on the recommended parameters of Warner et al. (2024). The label smoothing parameter $\alpha$ was set to 0.1. The seed was fixed at $42$ and a training batch size of $16$ and validation and test batch size of $64$ were used for all experiments. The learning rate was set based on a hyperparameter sweep over [$1$, $1.5$, $2$, $3$, $4$]×$\times 10^{-5}$. For ModernBERT the best-performing learning rate was $10^{-5}$. The classification threshold was selected to maximise the macro average F1 score on the validation set.
Appendix B Model collapse experiments
B.1 Dataset
All our model collapse experiments use the raw variant of the WikiText-2 dataset Merity et al. (2016). We train models on the training set, consisting of $36,718$ documents and evaluate on the test set of $4,358$ documents. The WikiText dataset was extracted from the Good or Featured article criteria specified by editors on Wikipedia and only covers the English language.
B.2 LLMs
The Generative Pre-trained Transformer 2 (GPT-2Radford et al. (2019) is a decoder-only transformer-based language model. GPT-2 demonstrated that large-scale language models could perform various language tasks without task-specific training. We use the base variant, which contains $117$M parameters. SmolLM2 Allal et al. (2025) is a family of compact and efficient language models developed by Hugging Face, available in three sizes: $135$M, $360$M, and $1.7$B parameters. We utilise the $360$M parameter variant in our experiments.
B.3 Hyperparameters
In our experiments, we conduct full fine-tuning using a learning rate of $5\!\times\!10^{-5}$, batch size of $8$ and a dropout rate of $0.1$. For the AdamW optimizer, we use $\beta_{1}$, $\beta_{2}$ and $\epsilon$ of $0.9$, $0.999$ and $10^{-8}$ respectively. Each model was trained for $1$ epoch with the hyperparameters fixed for all experiments. We conducted $10$ iterations of recursive training.
Appendix C Informed sampling
C.1 Bias term
To correct for the bias in the classifier’s confidence estimates toward predicting text as machine-generated, we introduce a bias factor $b$ based on the classification threshold:
$b=1+\frac{\text{threshold}}{1-\text{threshold}}.$ (S1)
This formulation is motivated by the odds ratio at the decision boundary, where $\frac{\text{threshold}}{1-\text{threshold}}$ represents the relative likelihood of a sample being classified as machine-generated versus human-generated at the threshold. By adding $1$, we ensure that $b\!\geq 1$, maintaining a minimum correction factor of 1. For a classifier with $\text{threshold}>0.5$, this formulation increases the selection probability of underrepresented samples, counteracting the bias introduced by the classifier’s skewed confidence distribution.
@@ C.1 Bias term
SmolLM2 baseline top-$k$5, 1, 0 $13.36$ $14.06$ $47.89$ $47.57$ $82.50$ $73.99$ $48.99$ $52.77$ $94.58$ $91.10$ $55.65$ $58.09$
.5, .5, .5 $13.36$ $14.00$ $47.89$ $47.58$ $82.46$ $76.24$ $48.66$ $52.44$ $93.68$ $95.13$ $55.24$ $57.88$
1, 1, 0 $13.36$ $13.49$ $47.88$ $47.85$ $82.86$ $77.80$ $48.68$ $51.36$ $94.76$ $89.06$ $55.43$ $57.27$
temperature .5, 1, 0 $13.36$ $14.07$ $47.89$ $47.70$ $83.42$ $66.25$ $45.57$ $52.77$ $95.37$ $81.27$ $52.72$ $57.67$
.5, .5, .5 $13.36$ $13.95$ $47.89$ $47.67$ $83.27$ $71.22$ $45.89$ $52.73$ $88.33$ $84.54$ $53.01$ $56.23$
1, 1, 0 $13.36$ $13.46$ $47.87$ $47.93$ $82.73$ $72.75$ $45.69$ $52.21$ $94.57$ $87.55$ $52.88$ $56.39$
ours top-$k$5, 1, 0 $13.36$ $13.56$ $47.88$ $47.76$ $82.21$ $81.77$ $49.04$ $49.91$ $94.96$ $93.19$ $55.49$ $56.64$
.5, .5, .5 $13.36$ $13.51$ $47.87$ $47.83$ $82.56$ $81.86$ $48.83$ $49.62$ $94.55$ $94.46$ $55.33$ $56.55$
1, 1, 0 $13.36$ $13.15$ $47.88$ $48.08$ $82.35$ $82.52$ $45.00$ $48.91$ $94.00$ $94.43$ $55.35$ $56.14$
temperature .5, 1, 0 $13.36$ $13.64$ $47.88$ $47.79$ $82.83$ $79.35$ $46.08$ $49.15$ $93.13$ $70.63$ $52.87$ $56.08$
.5, .5, .5 $13.36$ $13.62$ $47.88$ $47.80$ $83.31$ $79.62$ $45.64$ $49.34$ $93.07$ $96.14$ $52.79$ $55.68$
1, 1, 0 $13.36$ $13.26$ $47.87$ $48.06$ $83.24$ $80.87$ $45.94$ $48.50$ $93.93$ $78.36$ $52.64$ $54.90$
Table S1: Test performance (perplexity and accuracy) and data quality at generation 00 and generation $9$. Results are shown for top-$k$ decoding and sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ ($\uparrow$ / $\downarrow$: higher / lower is better).

### v5

@@ 7 Conclusion
This work investigates model collapse across three dimensions: model performance, data quality, and resemblance to human samples. Through our analysis, we found that the extent of model collapse and the effect on the data distribution is influenced by the decoding strategy. Truncating can lead to peaked distributions and repetitive models while pure sampling can result in high perplexity and verbose outputs with low resemblance to human data. Using the decoding strategies that resulted in the most extreme collapse, we evaluated the partially synthetic scenario, where human data is mixed into the training data. We designed a novel method to mitigate model collapse based on resampling the training distribution using the predictions of a machine-generated text detector. We have validated our method on two popular model variants (GPT-2 and SmolLM2) across a range of decoding strategies and model sizes, showing that we can prevent model collapse in all cases. When there is an equal ratio of human to synthetic samples in the training pool, our method results in improved model performance compared to training only on the human data.
@@
As in previous studies Shumailov et al. (2023); Dohmatob et al. (2024), we assess LLMs exclusively in a fine-tuning setting rather than pre-training from scratch. While pre-training experiments could provide deeper insights, the computational cost and complexity of training large-scale models from the ground up make such an approach impractical in our case. Nevertheless, given that model collapse has been primarily evaluated in LLMs from a fine-tuning setting, the conclusions made in this work still align with the current body of research.
In addition, our study focuses primarily on open-ended text generation tasks. While this is a crucial area for understanding model collapse, our findings may not fully generalise to other domains, such as structured prediction or code generation, where the impact of model collapse may manifest differently. Future work could explore whether our resampling method remains effective across these domains.
Finally, our method depends on the performance of the machine-generated text detector used to estimate the importance weights. Any inaccuracies or biases in the detector’s predictions directly affect the quality of the resampling process, potentially leading to suboptimal mitigation of model collapse. Since detector performance varies with domain, LLM architecture, and decoding strategy, the generalisability of our approach is closely tied to the detector’s robustness. The primary focus of this work, however, is not on optimising the detector itself, but on demonstrating that detector-based resampling can effectively mitigate model collapse. Future work could investigate improved methods of detection, such as using adaptive or ensemble-based detectors to improve reliability across data regimes and LLMs.
@@
Appendix A Recursive training
@@ A.1 Dataset
All our model collapse experiments use the raw variant of the WikiText-2 dataset Merity et al. (2016). We train models on the training set, consisting of $36{,}718$ documents, and evaluate on the test set of $4{,}358$ documents. The WikiText-2 dataset was extracted from the ‘Good’ or ‘Featured’ article criteria specified by editors on Wikipedia and only covers the English language.
A.2 LLMs
GPT-2 (Generative Pre-trained Transformer 2) Radford et al. (2019) is a decoder-only transformer-based language model. GPT-2 demonstrated that large-scale language models could perform various language tasks without task-specific training. We use the base variant, which contains $124$M parameters. SmolLM2 Allal et al. (2025) is a family of compact and efficient language models developed by Hugging Face, available in three sizes: $135$M, $360$M, and $1.7$B parameters. The majority of our experiments use the $360$M parameter variant unless specified otherwise.
A.3 Hyperparameters
In our experiments, we conduct full fine-tuning using a learning rate of $5\!\times\!10^{-5}$, batch size of $8$ and a dropout rate of $0.1$. For the AdamW optimizer, we set $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. Each model was trained for $1$ epoch with the hyperparameters fixed for all experiments. We conducted $10$ iterations of recursive training.
Appendix B Machine-generated text detection
B.1 Pre-trained models
B.2 Hyperparameters
Each model was fine-tuned for $5$ epochs. We select the best model based on the highest AUC on the validation set. Optimisation was performed using AdamW by setting $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=10^{-6}$, and the weight decay to $10^{-2}$. These parameters were chosen based on prior work Warner et al. (2024). The label smoothing parameter $\alpha$ was set to 0.1, the seed was fixed at $42$ and the training batch size to $16$. The learning rate was set based on a hyperparameter sweep over [$1$, $1.5$, $2$, $3$, $4$]×$\times 10^{-5}$. For ModernBERT the best-performing learning rate was $10^{-5}$. We implemented temperature scaling by learning the temperature parameter using L-BFGS optimisation on the validation set. This was run for $50$ iterations with a learning rate of $0.01$.
B.3 Dataset
@@
We trained and evaluated the machine-generated text detectors on the MAGE dataset Li et al. (2024), which is based on documents from $10$ domains: opinion statements (CMV & Yelp reviews dataset), news articles (XSum & TLDR dataset), question answering (ELI5), story generation (Reddit WritingPrompts & ROC), commonsense reasoning (HellaSwag), knowledge illustration (SQuAD) and Scientific writing (SciGen). The authors sampled $1{,}000$ texts from each domain (apart from opinion statements and news articles with $804$ and $777$ samples respectively) and generated text using $27$ LLMs from $7$ model families, which include OpenAI, LLaMA, GLM, FLAN-T5, OPT, BigScience and EleutherAI. For each human-written sample in the dataset, they generate a machine-generated version by providing the first $30$ tokens of human-written text as context to the LLM. In addition, for the OpenAI models, they implemented two other prompt strategies for relevant domains: ‘topical’ prompts such as an argument or news title and ‘specified’ prompts which contain information about the domain source. This results in $33{,}000$ ($=27{,}000+3\!\times\!2\!\times\!1{,}000$) machine-generated samples per source before processing and filtering. The authors split the dataset into train, validation and test splits in the ratio $80$:$10$:$10$. To mitigate data imbalance in the validation and test sets they sample additional human data from each data source. The resulting test set contains $28{,}741$ human and $28{,}078$ machine-generated samples ($49\%$ machine-generated). The training set, however, is $71\%$ machine-generated. The total dataset consists of $154{,}078$ human-written and $294{,}381$ machine-generated texts. In addition to the previously described test set, we also evaluate our detector on their more challenging test set containing text from four unseen domains (CNN/DailyMail, DialogSum, PubMedQA, IMDB) and generated by an unseen model (GPT-4). This out-of-distribution test set contains $762$ human and $800$ machine-generated samples.
When evaluating the ModernBERT model fine-tuned on MAGE on the SmolLM2 models, we observed a drop in performance compared to GPT-2, with large variability across decoding strategies and model size. For SmolLM2 $360$M the detector achieved a classification accuracy of .601.601.601.601 for top-$k$ decoding and .399.399.399.399 for temperature sampling. To ameliorate this, we finetuned a new ModernBERT model on a larger corpus, containing the MAGE dataset and a subset of the RAID dataset Dugan et al. (2024) for the SmolLM2 models. The RAID dataset is the largest machine-generated text detection dataset, includes text samples generated by $11$ LLMs with $4$ decoding strategies, and spans text across $8$ domains. Additionally, RAID includes $11$ types of adversarial attacks, such as homoglyph substitutions, number insertions, article deletions, and paraphrasing. We partitioned the dataset into training, validation, and test splits in the ratio $80$:$10$:$10$, ensuring no cross-contamination of text segments generated from the same source of text across splits. We balanced each split so that it contained an equal number of human and machine samples, stratified across model, decoding strategy, source domain, and adversarial attack (the whitespace and paragraph attacks were included). This resulted in balanced train, validation, and test splits comprising $128{,}352$, $16{,}044$, and $16{,}056$ samples, respectively.
Appendix C Informed sampling of training data
As we perform sampling with replacement, there is the risk of excessive duplication of high-weight samples. To account for this, we introduce a maximum resample count parameter, $r_{\text{max}}$, which limits the number of times any individual sample can be selected. This constraint ensures diversity in the resampled dataset and prevents a small subset of high-weight samples from dominating the training distribution. To further correct for the classifier’s bias toward labelling samples as machine-generated, we introduce a bias term $b\!\geq 1$ to adjust the weight distribution. This formulation increases the selection probability of samples likely to be human, counteracting the bias introduced by the classifier’s skewed confidence distribution. We select values for $r_{\text{max}}$ and $b$ by evaluating each model on the Wikitext-2 validation set after 1 generation of recursive training. The optimal hyperparameters for each model configuration are reported in Table S1.
Model $r_{\text{max}}$ $b$
GPT-2 10 10
SmolLM 135M 10 10
SmolLM 360M 5 10
SmolLM 1.7B 3 1
Table S1: Optimal hyperparameters for Sampling Importance Resampling across different model scales using top-$k$ decoding.
Refer to caption Refer to caption Refer to caption Refer to caption
GPT-2 SmolLM2
@@ Appendix C Informed sampling of training data
Figure S1: Perplexity and accuracy over generations 00 to $9$ of fully synthetic recursive training for varying mixing coefficients ($\alpha$, $\beta$, $\gamma$) using top-$k$ decoding.
Refer to caption Refer to caption Refer to caption Refer to caption
GPT-2 SmolLM2
Figure S2: Classification score distribution of the machine-generated text detector on the dataset from generation 00.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure S3: Perplexity distribution for machine-generated text of SmolLM2 at generations 00, $1$, and $9$ in the partially synthetic scenario $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$. Perplexity is evaluated using the model trained on human text ($p^{0}$).
@@
$1.7$B Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$360$M Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$135$M Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure S5: SmolLM2 model size variants ($1.7$B, $360$M, $135$M) under partially synthetic recursive training $\left(\alpha\!=\!1,\beta\!=\!1,\gamma\!=\!0\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
@@
SmolLM2 $350$M baseline top-$k$5, 1, 0 $13.96$ $14.69$ $47.58$ $47.30$ $82.59$ $73.95$ $52.68$ $54.53$ $91.58$ $78.42$ $55.31$ $58.10$
.5, .5, .5 $13.96$ $14.64$ $47.59$ $47.30$ $82.76$ $76.54$ $52.57$ $53.97$ $91.03$ $86.15$ $55.38$ $57.84$
1, 1, 0 $13.96$ $14.10$ $47.57$ $47.57$ $82.51$ $76.90$ $52.47$ $54.17$ $91.32$ $81.11$ $55.57$ $57.24$
temperature .5, 1, 0 $13.96$ $14.69$ $47.59$ $47.37$ $83.16$ $66.16$ $51.21$ $54.46$ $83.67$ $70.15$ $52.59$ $57.81$
.5, .5, .5 $13.96$ $14.60$ $47.58$ $47.36$ $83.74$ $72.03$ $51.22$ $54.15$ $86.94$ $81.27$ $52.80$ $56.86$
1, 1, 0 $13.96$ $14.06$ $47.58$ $47.64$ $83.31$ $73.11$ $51.16$ $53.96$ $84.55$ $76.27$ $52.74$ $55.85$
ours top-$k$5, 1, 0 $13.96$ $14.02$ $47.58$ $47.62$ $82.75$ $79.17$ $52.58$ $53.89$ $89.33$ $90.39$ $55.53$ $57.67$
.5, .5, .5 $13.96$ $14.00$ $47.59$ $47.59$ $82.52$ $79.79$ $52.64$ $53.74$ $88.69$ $87.92$ $55.39$ $57.41$
1, 1, 0 $13.96$ $13.67$ $47.60$ $47.83$ $82.93$ $81.67$ $52.74$ $53.23$ $85.09$ $89.72$ $55.53$ $56.90$
temperature .5, 1, 0 $13.96$ $14.11$ $47.58$ $47.63$ $83.45$ $75.84$ $51.15$ $53.67$ $87.80$ $84.82$ $53.03$ $56.74$
.5, .5, .5 $13.96$ $14.13$ $47.58$ $47.58$ $82.99$ $77.39$ $51.12$ $53.08$ $86.20$ $81.00$ $52.98$ $56.09$
1, 1, 0 $13.96$ $13.75$ $47.57$ $47.83$ $83.43$ $79.36$ $51.21$ $52.99$ $88.16$ $86.94$ $52.94$ $55.66$
SmolLM2 $135$M baseline top-$k$ 1, 1, 0 $19.02$ $19.35$ $44.08$ $44.04$ $79.08$ $70.99$ $53.08$ $54.87$ $85.18$ $72.73$ $57.34$ $59.49$
ours $19.02$ $18.54$ $44.10$ $44.37$ $79.15$ $78.06$ $52.90$ $53.69$ $82.92$ $82.24$ $57.35$ $59.04$
SmolLM2 $1.7$B baseline top-$k$ 1, 1, 0 $9.31$ $9.39$ $52.67$ $52.70$ $85.24$ $81.02$ $53.22$ $54.28$ $91.11$ $85.95$ $53.79$ $55.02$
ours $9.31$ $9.31$ $52.66$ $52.73$ $85.13$ $83.81$ $53.16$ $53.68$ $88.22$ $93.23$ $54.01$ $54.66$
Table S2: Test performance (perplexity and accuracy) and data quality at generation 00 and generation $9$. Results are shown for top-$k$ decoding and pure sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ ($\uparrow$ / $\downarrow$: higher / lower is better).